# Bundle-specific Tractogram Distribution Estimation Using Higher-order Streamline Differential Equation

### v1

@@ 2.2 Spatial continuity constraint equation for a tractogram
We assume that the diffusion displacement of water molecules in the same fiber maintains continuity. We use continuous incompressible fluid theory to describe the spatial continuity of the bundle by introducing the concept of divergence of the fiber flow on diffusion tensor vectorial field,
$divS=\frac{{\partial{v_{x}}}}{{\partial x}}+\frac{{\partial{v_{y}}}}{{\partial y}}+\frac{{\partial{v_{z}}}}{{\partial z}}.$ (10)
We assume that the fibers do not originate or terminate in the white matter, that is, $divS$ satisfies
$divS=0.$ (11)
@@
$\begin{array}[]{l}divS=\sum\limits_{i=0}^{n}{\sum\limits_{j=0}^{n-i}{\sum\limits_{k=0}^{n-i-j}{\left(\begin{array}[]{l}a_{ijk}^{x}i{x^{i-1}}{y^{j}}{z^{k}}+\\ a_{ijk}^{y}j{x^{i}}{y^{j-1}}{z^{k}}+\\ a_{ijk}^{z}k{x^{i}}{y^{j}}{z^{k-1}}\end{array}\right)}}}\\ {\rm{=}}{H_{n-1}}\cdot{C_{n-1}}(x,y,z)=0\end{array}$ (12)
@@
${H_{n-1}}={\left[\begin{array}[]{l}na_{n00}^{x}{\rm{+}}a_{(n-1){\rm{1}}0}^{y}{\rm{+}}a_{(n-1)01}^{z}\\ (n-1)a_{(n-1)10}^{x}{\rm{+2}}a_{{\rm{(n-2)2}}0}^{y}{\rm{+}}a_{(n-2)11}^{z}\\ (n-1)a_{(n-1)01}^{x}{\rm{+}}a_{(n-2)11}^{y}{\rm{+2}}a_{(n-2)02}^{z}\\ (n-1)a_{(n-1)00}^{x}{\rm{+}}a_{(n-2)10}^{y}{\rm{+}}a_{(n-2)01}^{z}\\ \cdot\cdot\cdot\\ {\rm{(}}i{\rm{+1)}}a_{(i+1)jk}^{x}+(j+1)a_{i(j+1)k}^{y}+\\ (k+1)a_{ij(k+1)}^{z}\\ \cdot\cdot\cdot\\ a_{101}^{x}{\rm{+}}a_{011}^{y}{\rm{+2}}a_{002}^{z}\\ a_{100}^{x}{\rm{+}}a_{010}^{y}{\rm{+}}a_{001}^{z}\end{array}\right]^{T}}_{\left(1,M\right)}$
@@ 2.3 Estimation of the BTD
With the definition of BTD, finding the optimal tractogram from ${\rho_{1}}$ to ${\rho_{2}}$ can be simplified as the estimation of coefficient matrix $A$. The coefficient matrix $A$ is estimated by minimizing the energy $P(S)$ in Eq. 4 with $p\left({S(t),\dot{S}(t)}\right)$ defined as:
@@
where $\Phi(v(x,y,z))$ is the probability that fiber trajectory is actually passes through the FOD at point $(x,y,z)$, which is derived from the orientation distribution functions (ODF) in each voxel in the volume. To achieve fiber spatial continuity, we add constrain Eq. 12 to Eq. 4, which yields the optimization model:
@@
$\mathop{\min}\limits_{A}E=\left\|{G-A\cdot C_{n}^{*}(X,Y,Z)}\right\|_{2}^{2}$
$s.t.{\rm{}}{H_{n-1}}\cdot C_{n-1}^{*}(X,Y,Z)=0$
$\begin{array}[]{l}\begin{split}C_{n}^{*}(X,Y,Z)=&[C_{n}({X_{1}},{Y_{1}},{Z_{1}}),C_{n}({X_{2}},{Y_{2}},{Z_{2}}),\\ &\cdot\cdot\cdot,C_{n}({X_{\gamma}},{Y_{\gamma}},{Z_{\gamma}})]_{\left({\sum\limits_{\Delta=0}^{n}{\frac{{(\Delta+1)*(\Delta+2)}}{2}},{\rm{}}\gamma}\right)}\end{split}\end{array}$ (15)
The estimation of the BTD is decomposed into two stages. First, the least squares method is used to solve Eq. 15 to obtain the coefficient matrix $A$. Second, we use the coefficient matrix $A$ to obtain the flow field vectors and use the fourth-order Runge-Kutta method [28] to solve the higher-order streamline differential equation Eq. 6. See Algorithm. 1 for a detailed implementation.
Input: A set of peak directions $G={[{g_{1}}{\rm{,}}{g_{2}}\cdot\cdot\cdot{g_{\gamma}}]_{\left({3,{\rm{}}\gamma}\right)}}$ , the center coordinate $(X,Y,Z)$ of each voxel in bundle mask $\Omega$, seed points $\left\{{{\beta_{i}},i=1,\cdots,\lambda}\right\}$, tractography step size $\Delta$.
@@ 3 Experiments
Tests and validations of our algorithm are based on four simulated datasets, namely, Hough data, Sine data [28], Circle data, and ISMRM 2015 Tractography Challenge data [1], and in vivo data of HCP [49]. Our BTD estimation is based on FODs estimated using constrained spherical deconvolution, which are included in the software package MRtrix [11]. The experimental results are evaluated by comparing BTD with deterministic FOD-based tracking (SD_Stream) [50], integration over fiber orientation distributions (iFOD2) [51], and unscented Kalman filter (UKF) algorithm [52]. For quantitative evaluation, we use Tractometer metrics, that is, valid connections (VC), bundle overlap (OL), and bundle overreach (OR) for Sine and Hough data [1, 53, 54]. For Circle data, we design a metric referred to as $Deviation$, which can be expressed as:
$Deviation=\sum\limits_{j=1}^{N}{\sum\limits_{i=1}^{{\lambda_{j}}}{(\sqrt{{{({x_{ij}}-{U_{c}})}^{2}}+{{({y_{ij}}-{V_{c}})}^{2}}}}}-{R_{j}})/\sum\limits_{j=1}^{N}{({\lambda_{j}})}$
@@ 3.3 HCP data
We used HCP #100307 subject to visually compare fiber tractograms from the proposed BTD algorithm and results from the tract orientation mapping (TOM) [57], parallel transport tractography (PTT) [33], fiber trajectory distribution (FTD) [34], unscented Kalman filter (UKF) algorithm [52], integration over fiber orientation distributions (iFOD2) [51], and deterministic FOD-based tracking (SD_Stream) [50]. In this paper, we use VC and OL to validate the proposed method, so all tractography methods were selected 3000 streamlines which were seeded from all voxels within the each tract start regions, and tract masks were used to filter the tractograms. The tractography specific parameters are as follows: i) FTD, iFOD2, and SD_Stream: maximum angle = $60^{0}$, step size = 0.3, cutoff = 0.1, minimum length = 75; ii) UKF: seedingFA = 0.06, stoppingFA = 0.05, stoppingThreshold = 0.06, Qm = 0.001, and Ql = 50. iii) PTT: the default parameters in [33], the difference is that the seeding regions are the tract start regions. iv) TOM: the default parameters in [57], the difference is that the seeding regions are the tract start regions, max_nr_fibers=3000, no streamline filtering by end mask.
In the absence of ground truth fibers for in vivo data, the relatively familiar corpus callosum (CC_4), CST, and arcuate fasciculus (AF) were selected for qualitative and quantitative evaluations. The AF is a neuronal pathway that connects Wernicke’s area and Broca’s area [58]. The CC_4 is minor forceps of the corpus callosum that connects bilateral frontal lobe [59]. These three tracts have the characteristics of long-range, twisting, and fanning characteristics, making them suitable to assess the algorithms.
@@
4 Discussion and Conclusion
@@
In this work, a novel bundle-specific tractography approach BTD is proposed, which integrates higher-order streamline differential equation to derive brain connectome between two regions in ‘cluster to cluster’ manner instead of ‘single to single’ manner. We parameterize fiber bundles using the BTD coefficients that are estimated by minimizing the energy on the diffusion tensor vector field by combining the priors. Experiments are performed on Hough, Sine, Circle data, and the ISMRM 2015 Tractography Challenge data and in vivo data of HCP for qualitative and quantitative evaluation. The horizontal comparisons show that with increasing order of the BTD, the numbers of valid fibers and overlapped regions will gradually increase. The best results are obtained with the 5th- or 6th-order BTD, and an order higher than six may cause overfitting. The comparisons with state-of-the-art methods show that the BTD can reconstruct complex fiber bundles, such as long-range, large twisting, and fanning tracts, and show better spatial consistency with fiber geometry, which is potentially useful for robust tractography.
Our method may be affected when FODs or peaks are inaccurate due to noise, artifacts, pathological cases, or poor-quality datasets. In addition, the BTD has difficulty tracking bundles that have received tumor compression. Fiber tracking using deep learning in tumor scenarios will be studied in future work.

### v2

@@ 2.2 Spatial continuity constraint equation for a tractogram
We assume that the diffusion displacement of water molecules in the same fiber maintains continuity. We use continuous incompressible fluid theory to describe the spatial continuity of the bundle by introducing the concept of divergence of the fiber flow on diffusion vectorial field. Let $div\Gamma$ denote the divergence of the fiber flow on diffusion vectorial field, which is formulated as
$div\Gamma=\frac{{\partial{v_{x}}}}{{\partial x}}+\frac{{\partial{v_{y}}}}{{\partial y}}+\frac{{\partial{v_{z}}}}{{\partial z}},$ (10)
where $v_{x}$, $v_{y}$, and $v_{z}$ represent vector value in $x$, $y$, and $z$ axis. We assume that the fibers do not originate or terminate in the white matter, that is, $div\Gamma$ satisfies
$div\Gamma=0.$ (11)
@@
$\begin{array}[]{l}div\Gamma=\sum\limits_{i=0}^{n}{\sum\limits_{j=0}^{n-i}{\sum\limits_{k=0}^{n-i-j}{\left(\begin{array}[]{l}a_{ijk}^{x}i{x^{i-1}}{y^{j}}{z^{k}}+\\ a_{ijk}^{y}j{x^{i}}{y^{j-1}}{z^{k}}+\\ a_{ijk}^{z}k{x^{i}}{y^{j}}{z^{k-1}}\end{array}\right){\rm{}}}}}\\ {\rm{=}}{H_{n-1}}\cdot{C_{n-1}}(x,y,z)=0\end{array}$ (12)
@@
${H_{n-1}}={\left[\begin{array}[]{l}na_{n00}^{x}{\rm{+}}a_{(n-1){\rm{1}}0}^{y}{\rm{+}}a_{(n-1)01}^{z}{\rm{}}\\ (n-1)a_{(n-1)10}^{x}{\rm{+2}}a_{{\rm{(n-2)2}}0}^{y}{\rm{+}}a_{(n-2)11}^{z}\\ (n-1)a_{(n-1)01}^{x}{\rm{+}}a_{(n-2)11}^{y}{\rm{+2}}a_{(n-2)02}^{z}\\ (n-1)a_{(n-1)00}^{x}{\rm{+}}a_{(n-2)10}^{y}{\rm{+}}a_{(n-2)01}^{z}\\ \cdot\cdot\cdot\\ {\rm{(}}i{\rm{+1)}}a_{(i+1)jk}^{x}+(j+1)a_{i(j+1)k}^{y}+\\ (k+1)a_{ij(k+1)}^{z}\\ \cdot\cdot\cdot\\ a_{101}^{x}{\rm{+}}a_{011}^{y}{\rm{+2}}a_{002}^{z}\\ a_{100}^{x}{\rm{+}}a_{010}^{y}{\rm{+}}a_{001}^{z}\end{array}\right]^{T}}_{\left(1,M\right)}$
@@ 2.3 Estimation of the BTD
With the definition of BTD, finding the optimal tractogram from ${\rho_{1}}$ to ${\rho_{2}}$ can be simplified as the estimation of coefficient matrix $A$. The coefficient matrix $A$ is estimated by minimizing the energy $\mathrm{E}(S)$ in Eq. 3 with $p\left({S(t),\dot{S}(t)}\right)$ defined as:
@@
where $\Phi(v(x,y,z))$ is the probability that fiber trajectory is actually passes through the FOD at point $(x,y,z)$, which is derived from the ODF in each voxel in the volume. To achieve fiber spatial continuity, we add constrain Eq. 12 to Eq. 3, which yields the optimization model:
@@
$\begin{array}[]{l}\mathop{\min}\limits_{A}E=\left\|{G-A\cdot C_{n}^{*}(X,Y,Z)}\right\|_{2}^{2}\\ s.t.{\rm{}}{H_{n-1}}\cdot C_{n-1}^{*}(X,Y,Z)=0\\ \begin{split}C_{n}^{*}(X,Y,Z)=&[C_{n}({X_{1}},{Y_{1}},{Z_{1}}),C_{n}({X_{2}},{Y_{2}},{Z_{2}}),\\ &\cdot\cdot\cdot,C_{n}({X_{\gamma}},{Y_{\gamma}},{Z_{\gamma}})]_{\left({\sum\limits_{\Delta=0}^{n}{\frac{{(\Delta+1)*(\Delta+2)}}{2}},{\rm{}}\gamma}\right)}\end{split}\end{array}$ (15)
The estimation of the BTD is decomposed into two stages. First, the least squares method is used to solve Eq. 15 to obtain the coefficient matrix $A$. Second, we use the coefficient matrix $A$ to obtain the flow field vectors and use the fourth-order Runge-Kutta method [30] to solve the higher-order streamline differential equation Eq. 7. See Algorithm. 1 for a detailed implementation. The process for generating the bundle masks in our experiments as follows: i) In vivo data: First, constrained spherical deconvolution (CSD) is applied to obtain the three principal FOD directions per voxel (Peaks) which is the input to the trained U-Nets [40]. Second, the optimal weights of the trained U-Nets are loaded to predict the given bundle mask and the given start/end region mask for each tract. ii) The simulated data: The bundle mask used were provided in the these public dataset.
Input: A set of peak directions $G={[{g_{1}}{\rm{,}}{g_{2}}\cdot\cdot\cdot{g_{\gamma}}]_{\left({3,{\rm{}}\gamma}\right)}}$ , the center coordinate $(X,Y,Z)$ of each voxel in bundle mask $\Omega=[{\omega_{1}},{\omega_{2}},\cdot\cdot\cdot,{\omega_{\gamma}}]$, seed points $\left\{{{\beta_{i}},i=1,\cdots,\lambda}\right\}$, tractography step size $\Delta$.
@@ 3 Experiments
Tests and validations of our algorithm are based on four simulated datasets, namely, Hough data, Sine data [30], Circle data, FiberCup data, and ISMRM 2015 Tractography Challenge data [1], and in vivo data of HCP [42]. Our BTD estimation is based on FODs estimated using constrained spherical deconvolution, which are included in the software package MRtrix [19]. The experimental results are evaluated by comparing BTD with deterministic FOD-based tracking (SD_Stream) [51], integration over fiber orientation distributions (iFOD2) [52], and unscented Kalman filter (UKF) algorithm [53]. For quantitative evaluation, we use Tractometer metrics, that is, valid connections (VC), bundle overlap (OL), and bundle overreach (OR) for Sine and Hough data [1, 54, 55]. For Circle data, we design a metric referred to as $Deviation$, which can be expressed as:
$Deviation=\sum\limits_{j=1}^{N}{\sum\limits_{i=1}^{{\lambda_{j}}}{(\sqrt{{{({x_{ij}}-{U_{c}})}^{2}}+{{({y_{ij}}-{V_{c}})}^{2}}}}}-{R_{j}})/\sum\limits_{j=1}^{N}{({\lambda_{j}})},$ (16)
@@ 3.3 HCP data
We used HCP #100307 subject to visually compare fiber tractograms from the proposed BTD algorithm and results from the tract orientation mapping (TOM) [40], parallel transport tractography (PTT) [29], fiber trajectory distribution (FTD) [49], unscented Kalman filter (UKF) algorithm [53], integration over fiber orientation distributions (iFOD2) [52], and deterministic FOD-based tracking (SD_Stream) [51]. In this paper, we use VC and OL to validate the proposed method, so all tractography methods were selected 3000 streamlines which were seeded from all voxels within the each tract start regions, and tract masks were used to filter the tractograms. The tractography specific parameters are as follows: i) FTD, iFOD2, and SD_Stream: maximum angle = $60^{0}$, step size = 0.3, cutoff = 0.1, minimum length = 75; ii) UKF: seedingFA = 0.06, stoppingFA = 0.05, stoppingThreshold = 0.06, Qm = 0.001, and Ql = 50. iii) PTT: the default parameters in [29], the difference is that the seeding regions are the tract start regions. iv) TOM: the default parameters in [40], the difference is that the seeding regions are the tract start regions, max_nr_fibers = 3000, no streamline filtering by end mask. For the HCP data, the start and end point of our proposed BTD use the start and end region segmented by the TractSeg [59] method. To ensure the fairness of the experiment, all competing methods use the same start point, end point and the bundle mask.
In the absence of ground truth fibers for in vivo data, the relatively familiar corpus callosum (CC_4), CST, and arcuate fasciculus (AF) were selected for qualitative and quantitative evaluations. The AF is a neuronal pathway that connects Wernicke’s area and Broca’s area [60]. The CC_4 is minor forceps of the corpus callosum that connects bilateral frontal lobe. These three tracts have the characteristics of long-range, twisting, and fanning characteristics, making them suitable to assess the algorithms.
@@
4 Discussion
In this paper, we define a bundle-specific tractogram distribution (BTD) function based on any higher-order streamline differential equation in the measured diffusion vectorial field. The results across the six different datasets show that the proposed BTD is capable of reconstructing complex fiber bundles with long distances, large twists, and fan-shaped bundles, and shows better spatial consistency with the fiber geometry. The order is an important parameter in our proposed method, so we give the effect of different orders on the proposed BTD and from the experimental results we can see that the 5th- or 6th-order BTD is the better parameter. Additionally, we also tested the computational speed required by the proposed algorithm and the results show that our method is reasonable time. Last but not least, our approach can accomplish fiber reconstruction without the need for expert ROIs placement, which also lowers inter-operator bias.
In our previous works [49], we designed the fiber trajectory distribution (FTD) function defined on the neighborhood voxels by using a ternary quadratic polynomial-based streamline differential equation, which can reveal continuous asymmetric fiber trajectory. However, the FTD is still a local reconstruction of the fiber trajectory at the neighborhood voxel level. Different from the FTD function, the proposed BTD formulates the tractography problem in the Riemannian manifold and integrates higher-order streamline differential equations to derive bundle-specific tractography. The BTD does not just is to fit a divergence-free vector field, but is simplified as the estimation of BTD coefficients by minimizing the energy optimization model, and is used to characterize the relations between BTD and diffusion tensor vector under the prior guidance by introducing the tractogram bundle information to provide anatomic priors. There is a reason why the proposed method provides such a good performance. The proposed BTD is to convert the fiber bundle connectivity relations modeled in terms of a voxelwise or local into the flow of the overall fluid on the Riemannian manifold. In the BTD function, the tangent vector at each point of the fiber path equals to field vector of the diffusion tensor, which is consistent with the definition of fiber tracking [1, 6]. The BTD function estimation algorithm takes pre-computed peak directions, voxel centroids, seed points, and tractography steps as input, and results in a bundle-specific tractogram.
In this paper, we define a new BTD function to describe the tract-wise tractogram. In order to ensure the fairness of the experiment, the bundle mask used in all our experiments was kept the same as the TOM method [59] [40]. For a given bundle mask, the proposed BTD function uses it initially for localization, then aims for global optimization within this localized region. Since all the comparison methods have slightly different approaches and requirements, for a fair comparison, we impose some restrictions on the other methods (TOM [40], PTT [29], FTD [49], UKF [53], iFOD2 [52], and SD_Stream [51]) when validating on vivo data from the Human Connectome Project (HCP) data, with the detailed parameters in Section 3.3. In this paper, we use VC and OL to validate the proposed method, so we set all comparison experiments to set the same seeds in the same start region, using the mask as a regional restriction for fiber tracking to compare the results. There is experimental evidence [1] [61] that complete seeding on the mask and the ROIs have an impact on the performance of fiber tracking. Therefore, there are some differences between the TOM we reproduced and the original results.
Tractography algorithms that offer better tools for examining brain topographic patterns are beneficial for both clinical and fundamental science applications. By integrating higher-order streamline differential equations, the proposed BTD poses the tractography issue in Riemannian manifolds and results in the derivation of bundle-specific tractography. We believe that the proposed BTD opens up new possibilities for structural brain mapping. We think that as neural network techniques progress, BTD can be used to provide a more accurate trajectory flow for any traction approach, not just for direct fiber bundle trajectory tracking but also as a direction via the fiber bundle region as a priori.
The following are some possible limitations of this study as well as some potential future research to solve these problems. Firstly, errors in the FOD or peaks of a single voxel will not affect our proposed method, but it will be affected when there are numerous errors due to noise, artifacts, pathological conditions, or low-quality datasets. More work is needed to directly utilize the original diffusion MRI signals instead of peaks in the fiber orientation distribution to fit the flow field direction. Secondly, although an exact segmentation of fiber bundles is not necessary, the presence of multiple fiber bundles within a single mask will affect the proposed method. Thus, the future study will therefore focus on figuring out how to model the ideal transport on the flow field and how to address the issue of false positive fibers showing up during the tracking process.
5 Conclusion
@@
In this work, a novel bundle-specific tractography approach BTD is proposed, which integrates higher-order streamline differential equation to derive brain connectome between two regions. We parameterize fiber bundles using the BTD coefficients that are estimated by minimizing the energy on the diffusion vectorial field by combining the priors. Experiments are performed on Hough, Sine, Circle data, FiberCup data, the ISMRM 2015 Tractography Challenge data, and in vivo data of HCP for qualitative and quantitative evaluation. The horizontal comparisons show that with increasing order of the BTD, the numbers of valid fibers and overlapped regions will gradually increase. The best results are obtained with the 5th- or 6th-order BTD, and an order higher than six may cause overfitting. The comparisons with state-of-the-art methods show that the BTD can reconstruct complex fiber bundles, such as long-range, large twisting, and fanning tracts, and show better spatial consistency with fiber geometry, which is potentially useful for robust tractography.
Data and code availability
The FiberCup data and the ISMRM 2015 Tractography Challenge data are available at  https://tractometer.org/. The Human Connectome Project dataset is available at  https://db.humanconnectome.org. The code of our method is openly available at https://github.com/IPIS-XieLei/BTD-Tractography.
Declarations of Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work was sponsored in part by the Zhejiang Province Science and Technology Innovation Leading Talent Program (No. 2021R52004), National Natural Science Foundation of China (Grant No.U22A2040, U23A20334, and 62303413) and Natural Science Foundation of Zhejiang Province (Grant No. LQ23F030017).